# On Deterministically Finding an Element of High Order Modulo a Composite

### v2

@@ 1. Introduction
When $N=p$ is prime, $\mathbb{Z}_{p}^{*}$ is cyclic, and deterministically finding a generator for this group is another very interesting open problem. To the best of our knowledge, the best known deterministic algorithm for this problem follows from using the factorization algorithms of [Har21, HH22b] to obtain a factorization of $p-1$ as $p-1=\prod_{i=1}^{m}q_{i}^{e_{i}}$, and then, using brute-force search, finding $q_{i}$-th non-residue $a_{i}$ modulo $p$ for every $q_{i}$, and outputting $\prod_{i=1}^{m}a_{i}^{(p-1)/q_{i}^{e_{i}}}$. Classical upper bounds on the magnitude of the least $q$-th non-residues [Bur57, Wan64, Nor71, Tre15] imply that the time complexity of the algorithm is dominated by the factorization step, so the total running time is $p^{1/5+o(1)}$. In [Gro15], Grossman presents a pseudo-deterministic algorithm for finding a generator modulo $p$ in time $\exp(\tilde{O}(\log p))$ (a pseudo-deterministic algorithm is a randomized algorithm that on every input, is guaranteed to output some canonical output with high probability).
@@ 1.2. Proof Technique
In a high level overview, one can think of Hittmeir’s algorithm as performing a brute force search for an element of high order, starting from $a=2,3,\ldots$ Given a specific element $a\in\mathbb{Z}_{N}^{*}$, one can deterministically check in time $O(\sqrt{D})\cdot\log^{O(1)}N$ whether the order of $a$ is at least $D$, using a simple variant on Shanks’s [Sha71] famous Baby-Step Giant-Step method for solving discrete logarithm: indeed, one can compute (and store) the list of elements $a^{j}$ for $10\leq j\leq\lceil\sqrt{D}\rceil-$ 1 (these are the “baby steps”) and the list of elements $a^{-i\lceil\sqrt{D}\rceil}$ for $10\leq i\leq\lceil\sqrt{D}\rceil-$ 1 (these are the “giant steps”), sort the lists and check for a collision in linear time (in the lengths of the lists). If indeed $a^{k}\equiv 1\bmod N$ for some $k\leq D$, then by writing $k=i\lceil\sqrt{D}\rceil+j$ for $10\leq i,j\leq\lceil\sqrt{D}\rceil-$ 1, we are guaranteed to find a collision, and in fact in this case we also know the exact order of $a$. If a collision is not found, then the order of $a$ is greater than $D$ (see Theorem 2.1 for a small improvement that shaves off a $\sqrt{\log\log D}$ factor).
@@
However, we can even improve the parameters a bit further. Since the other steps of our algorithm already run in time $D^{1/2+o(1)}$, we can allow the factorization step to run for longer than polynomial time. To that end, we adapt the technique of Coppersmith et al. [CHN08] to a recent deterministic algorithm of Harvey and Hittmeir [HH22a] for finding all $r$-th power divisors of an integer $N$. Their algorithm also builds on Coppersmith’s method [Cop97], which they adapt to exploit the fact that $N$ has an $r$-th power divisors, enabling them to find all such divisors in time $N^{1/4r}$.
@@
To search a given interval of the form $[P-H,P+H]$, Harvey and Hittmeir construct the system of polynomials
@@
Harvey and Hittmeir then cover the entire search range $[1,N^{1/r}]$ by such shorter intervals, and since the above procedure runs in polynomial time, the dominant contribution to the algorithm’s running time arises from the number of subintervals $[P-H,P+H]$ that must be processed.
@@
In Harvey and Hittmeir’s algorithm, [HH22a], the basic polynomial of interest is $f(x)=x+P$, where $P$ is the center of the current search interval $[P-H,P+H]$. The two important properties of this polynomial are:
@@
We remark that our algorithm, with straightforward modifications, works equally well for finding divisors in other residue classes modulo $s$ (i.e., divisors that equal $t$ modulo $s$ for some $10\leq t\leq s-$ 1). Since this is less useful in our context, we omit the details.
@@ 2. Preliminaries
In this section, we cite and prove some preliminary results. Sections 2.1, 2.2, and 2.3 summarize important subroutines we use in our algorithms. In Section 2.4 we prove some new results that we rely on for our improvement.
@@ 2.4. Number of Consecutive Roots
Let $k,\ell$ be positive integers such that $k\leq\frac{2}{\sqrt{5}}\cdot\sqrt{\ell}+1$ and $p$ a prime such that $p>2\ell$. Let $S=\{(x+a_{i})^{\ell}-\theta_{i}\}_{1\leq i\leq k}$ be system of univariate polynomials, where $\theta_{i},a_{i}\in\mathbb{F}_{p}$ and $a_{i}$’s are distinct. Then, the number of common roots of the polynomials in $S$ is at most $32\ell/(k-1)+$ 3.
@@
We note that 2.5 assumes that $N$ has a prime factor which is greater than $2\ell$. The reason we will be able to get away with this assumption in our main algorithm is that if $N$ has small prime factors, we can use Theorem 2.2 to find them quickly.
@@ 3. Factoring $N$ Given the Residue of a Prime Divisor
(3.1) $s\geq N^{\alpha}\text{ for some }0<\alpha<\tfrac{1}{2},\quad\gcd(N,s)=1.$
@@
Let $N,s\in\mathbb{N}$ with $\gcd(N,s)=1$. Let $s^{\prime}$ denote the unique integer in $[1,N-1]$ which is equivalent to the inverse of $s$ modulo $N$ [^2] We choose to use the notation $s^{\prime}$ rather than $s^{-1}$ in order to clearly distinguish between multiplication by $s^{\prime}$ and division by $s$: these operations are equivalent modulo $N$ but distinct over the integers.
Given $H<P\in\mathbb{N}$, suppose $p_{0}=sx_{0}+1$ is a divisor of $N$ that lies in the interval $[P-H,P+H]$. Let $\tilde{P}$ be the unique integer in $[0,s-1]$ which is equivalent to $P$ modulo $s$, and define $g(x)\in\mathbb{Z}[x]$ as
$g(x)\coloneq x+s^{\prime}+s^{\prime}(P-\tilde{P}).$
where every coefficient is taken modulo $N$ (so that every coefficient of $g$ is an integer in $[0,N-1]$). Further, let $x^{\prime}\coloneq x_{0}-\frac{P-\tilde{P}}{s}$. Then:
@@
Since $P\equiv\tilde{P}\pmod{s}$, we have that $s\mid(P-\tilde{P})$, meaning that $(P-\tilde{P})/s$ is an integer, and thus $x^{\prime}$ is also an integer. We observe that $(P-\tilde{P})/s\equiv s^{\prime}(P-\tilde{P})\pmod{N}$, and that $s^{\prime}p_{0}=s^{\prime}(sx_{0}+1)\equiv x_{0}+s^{\prime}\pmod{N}$. Therefore,
$g(x^{\prime})=x_{0}-(P-\tilde{P})/s+s^{\prime}+s^{\prime}(P-\tilde{P})$
@@
3.3 is used in order to search for prime divisors of $N$ (that equal 1 modulo $s$) in small intervals of the form $[P-H,P+H]$.
@@
The following theorem states that assuming $H$ is small enough, we may efficiently search intervals of the form $[P-H,P+H]$ for prime divisors that equal $1$ modulo $s$. In order to make the technical parameters easier to digest, the reader may assume that the parameters $d$ and $m$ are roughly logarithmic in $N$ (this will be true in the final setting of the parameters). The theorem is analogous to Theorem 2.1 in [HH22a], and the proof is very similar.
@@
Let $N,s\in\mathbb{N}$ be integers such that $s\geq N^{\alpha}$ for some $20<\alpha<1/$ 2, and $\gcd(N,s)=1$. Let $r\leq\log N$ be a positive integer, $d\in\mathbb{N}$ and $m\leq d/r$.
@@
there exists a deterministic algorithm that finds all integers $p$ such that $p^{r}\mid N$ and $p\equiv 1\pmod{s}$ in the interval $[P-H,P+H]$, that runs in time
@@
We can now run the LLL basis reduction algorithm [LLL82] on the lattice spanned by the vectors $\bar{v}_{i}$ for $0\leq i<d$. This algorithm returns a vector $\bar{w}$ which is an integral linear combination of the $\bar{v}_{i}$’s, such that
@@
The running time of the LLL basis reduction algorithm is polynomial in the dimension of the lattice $d$ and the logarithm of the norm of the basis vectors. Since the entries of $\bar{v}_{i}$ are bounded by $N^{O(d)}$, we can find a non-zero vector $\bar{w}$ in the lattice such that
@@
which is a linear combination of the polynomials $f_{i}(x)$ for $0\leq i<d$. Denoting $h=\sum_{i=0}^{d-1}h_{i}x^{i}$, we have that $h_{j}=w_{j}/G^{j}$. It follows that
@@
Next, as in [HH22a, Section 3], we consider finding $r$-power divisors $T\leq p^{r}\leq T^{\prime}$ in larger intervals $[T,T^{\prime}]$. The method would be to apply Theorem 3.5 to subintervals of the form $[P-H,P+H]$ and cover the entire interval $[T,T^{\prime}]$. As mentioned in [HH22a], it turns out that the running time is mainly determined by the number of subintervals, so we aim to maximize the size of $H$ while ensuring that the condition in Theorem 3.5 holds, mainly condition (3.2). For that, we require $G\leq\tilde{G}$ for
@@
The next proposition addresses the detection of $r$-th power divisors within the interval $[T,T^{\prime}]$, and is analogous to [HH22a, Proposition 3.2].
@@
the algorithm outputs a list of all integers $p$ within the interval $[T,T^{\prime}]$ that satisfy both $p^{r}\mid N$ and $p\equiv 1\pmod{s}$.
@@
[…]
@@
Let $H=(G-1)s$ so that $G=\lceil H/s\rceil+1$. Then $H\geq\frac{G}{2}s$, so that
@@
We now apply Theorem 3.5 with $N,r,d,m,G$ and $H$ as above and as set in Lemma 3.6, with $P=T+H,P=T+3H$ and so on, to cover the entire interval $[T,T^{\prime}]$.
@@
$\left\lceil\frac{T^{\prime}-T}{2H}\right\rceil\leq\frac{T^{\prime}-T}{\frac{s}{12}N^{\theta^{2}/2r}}+1=O\left(\frac{T^{\prime}-T}{T}N^{\theta(1-\theta)/r-\alpha}\right)$
@@
The proof of closely follows the argument of Theorem 1.1 in [HH22a].
@@
Letting $2^{k}=4^{\sqrt{\log N/r}}$, we now apply Proposition 3.7 successively on the intervals $[2^{k},2^{k+1}]$, $[2^{k+1},2^{k+2}]$ and so on, until reaching $N^{1/r}$. Since $k\geq 2\sqrt{\log N/r}$, the assumption of Proposition 3.7 is satisfied, and for each such interval, $\frac{T^{\prime}-T}{T}\leq 2$ and since $\theta\in[0,1]$, $\theta(1-\theta)\leq\frac{1}{4}$.
The number of intervals $[2^{k},2^{k+1}]$ is $O(\log N)$, and therefore the total running time is as stated in the theorem. ∎
@@ 4. The Main Algorithm
Algorithm 4.1 Finding an Element with Large Order or a Nontrivial Factor of $N$
@@
9:         Return $a$ as a nontrivial factor of $N$ or, if $a=N$, “$N$ is prime”.
@@
16:              Return $\gcd(N,a^{m_{i}/q}-1)$ as a nontrivial factor of $N$
@@
20:              Return $\gcd(N,M_{i+1})$ as a nontrivial factor of $N$
@@
By applying 2.5, the first while loop in step 6 takes $O(\sqrt{M_{i}})$ iterations (rather than the $O(M_{i})$ given by [Hit18]). This allows us to modify the condition in step 18 to $M_{i}\geq D$ instead of $M_{i}\geq\sqrt{D}$ without impacting the overall running time.
@@
By applying 2.5 with $\ell\coloneqq D$, we observe that the while loop in step 6 executes at most $O(\sqrt{D})$ iterations. Note that the condition that $N$ has a prime factor $p$ such that $p>2D$ is satisfied, otherwise a factor of $N$ would be found in step 1.
@@
[…]

### v3

@@ 1. Introduction
When $N=p$ is prime, $\mathbb{Z}_{p}^{*}$ is cyclic, and deterministically finding a generator for this group is another very interesting open problem. To the best of our knowledge, the best known deterministic algorithm for this problem follows from using the factorization algorithms of [Har21, HH22b] to obtain a factorization of p1p-1 as $p-1=\prod_{i=1}^{m}q_{i}^{e_{i}}$, and then, using brute-force search, finding $q_{i}$-th non-residue $a_{i}$ modulo $p$ for every $q_{i}$, and outputting $\prod_{i=1}^{m}a_{i}^{(p-1)/q_{i}^{e_{i}}}$. Classical upper bounds on the magnitude of the least $q$-th non-residues [Bur57, Wan64, Nor71, Tre15] imply that the time complexity of the algorithm is dominated by the factorization step, so the total running time is $p^{1/5+o(1)}$. In [Gro15], Grossman presents a pseudo-deterministic algorithm for finding a generator modulo $p$ in time $\exp(\tilde{O}(\log p))$ (a pseudo-deterministic algorithm is a randomized algorithm that on every input, is guaranteed to output some canonical output with high probability).
@@ 1.2. Proof Technique
In a high level overview, one can think of Hittmeir’s algorithm as performing a brute force search for an element of high order, starting from $a=2,3,\ldots$ Given a specific element $a\in\mathbb{Z}_{N}^{*}$, one can deterministically check in time $O(\sqrt{D})\cdot\log^{O(1)}N$ whether the order of $a$ is at least $D$, using a simple variant on Shanks’s [Sha71] famous Baby-Step Giant-Step method for solving discrete logarithm: indeed, one can compute (and store) the list of elements $a^{j}$ for $0\leq j\leq\lceil\sqrt{D}\rceil-1$ (these are the “baby steps”) and the list of elements $a^{-i\lceil\sqrt{D}\rceil}$ for $0\leq i\leq\lceil\sqrt{D}\rceil-1$ (these are the “giant steps”), sort the lists and check for a collision in linear time (in the lengths of the lists). If indeed $a^{k}\equiv 1\bmod N$ for some $k\leq D$, then by writing $k=i\lceil\sqrt{D}\rceil+j$ for $0\leq i,j\leq\lceil\sqrt{D}\rceil-1$, we are guaranteed to find a collision, and in fact in this case we also know the exact order of $a$. If a collision is not found, then the order of $a$ is greater than $D$ (see Theorem 2.1 for a small improvement that shaves off a $\sqrt{\log\log D}$ factor).
@@
However, we can even improve the parameters a bit further. Since the other steps of our algorithm already run in time $D^{1/2+o(1)}$, we can allow the factorization step to run for longer than polynomial time. To that end, we adapt the technique of Coppersmith et al. [CHN08] to a recent deterministic algorithm of Harvey and Hittmeir [HH22a] for finding all $r$-th power divisors of an integer $N$. Their algorithm also builds on Coppersmith’s method [Cop97], which they adapt to exploit the fact that $N$ has an $r$-th power divisor, enabling them to find all such divisors in time $N^{1/4r}$.
@@
To search a given interval of the form [PH,P+H][P-H,P+H], Harvey and Hittmeir construct the system of polynomials
@@
Harvey and Hittmeir then cover the entire search range $1,N^{1/r}]$ by such shorter intervals, and since the above procedure runs in polynomial time, the dominant contribution to the algorithm’s running time arises from the number of subintervals [PH,P+H][P-H,P+H] that must be processed.
@@
In Harvey and Hittmeir’s algorithm, [HH22a], the basic polynomial of interest is $f(x)=x+P$, where $P$ is the center of the current search interval [PH,P+H][P-H,P+H]. The two important properties of this polynomial are:
@@
We remark that our algorithm, with straightforward modifications, works equally well for finding divisors in other residue classes modulo $s$ (i.e., divisors that equal $t$ modulo $s$ for some $0\leq t\leq s-1$). Since this is less useful in our context, we omit the details.
@@ 2. Preliminaries
In this section, we cite and prove some preliminary results. Sections 2.1, 2.2, 2.3 and 2.4 summarize important subroutines we use in our algorithms. In Section 2.5 we prove some new results that we rely on for our improvement.
@@ 2.3. Order of All Prime Divisors
2.4. Lattice Basis Reduction
An important tool that we use in Section 3 is the famous LLL [LLL82] lattice basis reduction algorithm.
Lemma 2.4 ([LLL82]).
Let $d\geq 1$ and $B\geq 2$. Given a set of linearly independent vectors $v_{0},\dots,v_{d-1}\in\mathbb{Z}^{d}$ such that $\lVert v_{i}\rVert\leq B$ for all $i$, there exists a deterministic algorithm that runs in time polynomial in $d$ and $\log B$ and returns a nonzero vector $w$ in the lattice $L\coloneqq\operatorname{span}_{\mathbb{Z}}(v_{0},\dots,v_{d-1})$ such that
$\lVert w\rVert\leq 2^{(d-1)/4}\left(\det L\right)^{1/d}$
@@ 2.5. Number of Consecutive Roots
Let $k,\ell$ be positive integers such that $k\leq\frac{2}{\sqrt{5}}\cdot\sqrt{\ell}+1$ and $p$ be a prime such that $p>2\ell$. Let $S=\{(x+a_{i})^{\ell}-\theta_{i}\}_{1\leq i\leq k}$ be a system of univariate polynomials, where $\theta_{i},a_{i}\in\mathbb{F}_{p}$ and the $a_{i}$’s are distinct. Then, the number of common roots of the polynomials in $S$ is at most $2\ell/(k-1)+3$.
@@
We note that 2.6 assumes that $N$ has a prime factor which is greater than $2\ell$. The reason we will be able to get away with this assumption in our main algorithm is that if $N$ has small prime factors, we can use Theorem 2.2 to find them quickly.
@@ 3. Factoring $N$ Given the Residue of a Prime Divisor
(3.1) $s\geq N^{\alpha}\text{ for some }0<\alpha\leq\tfrac{1}{4r},\quad\gcd(N,s)=1.$
@@
Let $N,s\in\mathbb{N}$ with $\gcd(N,s)=1$. Let $s^{\prime}$ denote the unique integer in [1,N1][1,N-1] which is equivalent to the inverse of $s$ modulo NN. [^2] We choose to use the notation $s^{\prime}$ rather than $s^{-1}$ in order to clearly distinguish between multiplication by $s^{\prime}$ and division by $s$: these operations are equivalent modulo $N$ but distinct over the integers.
Given $H<P\in\mathbb{N}$, suppose $p_{0}=sx_{0}+1$ is a divisor of $N$ that lies in the interval [PH,P+H][P-H,P+H]. Let $\tilde{P}$ be the unique integer in [0,s1][0,s-1] which is equivalent to $P$ modulo $s$, and define $g(x)\in\mathbb{Z}[x]$ as
[…]
where every coefficient is taken modulo $N$ (so that every coefficient of $g$ is an integer in [0,N1][0,N-1]). Further, let $x^{\prime}\coloneq x_{0}-\frac{P-\tilde{P}}{s}$. Then:
@@
[…]
@@
3.3 is used in order to search for prime divisors of $N$ (that equal 1 modulo $s$) in small intervals of the form [PH,P+H][P-H,P+H].
@@
The following theorem states that assuming $H$ is small enough, we may efficiently search intervals of the form [PH,P+H][P-H,P+H] for prime divisors that equal $1$ modulo $s$. In order to make the technical parameters easier to digest, the reader may assume that the parameters $d$ and $m$ are roughly logarithmic in $N$ (this will be true in the final setting of the parameters). The theorem is analogous to Theorem 2.1 in [HH22a], and the proof is very similar.
@@
Let $N,s\in\mathbb{N}$ be integers such that $s\geq N^{\alpha}$ for some $0<\alpha<1/2$, and $\gcd(N,s)=1$. Let $r\leq\log N$ be a positive integer, $d\in\mathbb{N}$ and $m\leq d/r$.
@@
there exists a deterministic algorithm that finds all integers $p$ such that $p^{r}\mid N$ and $p\equiv 1\pmod{s}$ in the interval [PH,P+H][P-H,P+H], that runs in time
@@
We can now run the LLL [LLL82] basis reduction algorithm from Lemma 2.4 on the lattice spanned by the vectors $\bar{v}_{i}$ for $0\leq i<d$. This algorithm returns a vector $\bar{w}$ which is an integral linear combination of the $\bar{v}_{i}$’s, such that
@@
The running time of the LLL basis reduction algorithm is polynomial in the dimension of the lattice $d$ and the logarithm of the norm of the basis vectors (see Section 2.4). Since the entries of $\bar{v}_{i}$ are bounded by $N^{O(d)}$, we can find a non-zero vector $\bar{w}$ in the lattice such that
@@
which is a linear combination of the polynomials $f_{i}(x)$ for $0\leq i<d$. Denoting $h=\sum_{i=0}^{d-1}h_{i}x^{i}$, we have that $h_{j}=w_{j}/G^{j}$ (note that $h_{j}$ is an integer since $w_{j}$ is a linear combination of the integers $a_{i,j}$’s multiplied by $G^{j}$). It follows that
@@
Next, as in [HH22a, Section 3], we consider finding $r$-power divisors $T\leq p^{r}\leq T^{\prime}$ in larger intervals $T,T^{\prime}]$. The method would be to apply Theorem 3.5 to subintervals of the form [PH,P+H][P-H,P+H] and cover the entire interval $T,T^{\prime}]$. As mentioned in [HH22a], it turns out that the running time is mainly determined by the number of subintervals, so we aim to maximize the size of $H$ while ensuring that the condition in Theorem 3.5 holds, mainly condition (3.2). For that, we require $G\leq\tilde{G}$ for
@@
The next proposition addresses the detection of $r$-th power divisors within the interval $T,T^{\prime}]$, and is analogous to [HH22a, Proposition 3.2].
@@
the algorithm outputs a list of all integers $p$ within the interval $T,T^{\prime}]$ that satisfy both $p^{r}\mid N$ and $p\equiv 1\pmod{s}$.
@@
$\left(\frac{T^{\prime}-T}{T}\cdot N^{\theta(1-\theta)/r-\alpha}\right)\cdot{\mathsf{polylog}}(N)$
@@
Let H=(G1)sH=(G-1)s so that $G=\lceil H/s\rceil+1$. Then $H\geq\frac{G}{2}s$, so that
@@
We now apply Theorem 3.5 with $N,r,d,m,G$ and $H$ as above and as set in Lemma 3.6, with $P=T+H,P=T+3H$ and so on, to cover the entire interval $T,T^{\prime}]$.
@@
$\left\lceil\frac{T^{\prime}-T}{2H}\right\rceil\leq\frac{T^{\prime}-T}{\frac{s}{12}N^{\theta^{2}/r}}+1=O\left(\frac{T^{\prime}-T}{T}N^{\theta(1-\theta)/r-\alpha}\right)$
@@
The proof closely follows the argument of Theorem 1.1 in [HH22a].
@@
Letting $2^{k}=4^{\sqrt{\log N/r}}$, we now apply Proposition 3.7 successively on the intervals $2^{k},2^{k+1}]$, $2^{k+1},2^{k+2}]$ and so on, until reaching $N^{1/r}$. Since $k\geq 2\sqrt{\log N/r}$, the assumption of Proposition 3.7 is satisfied, and for each such interval, $\frac{T^{\prime}-T}{T}\leq 2$ and since $\theta\in[0,1]$, $\theta(1-\theta)\leq\frac{1}{4}$.
The number of intervals $2^{k},2^{k+1}]$ is $O(\log N)$, and therefore the total running time is as stated in the theorem. ∎
[…]
@@ 4. The Main Algorithm
Algorithm 4.1 Finding an Element with Large Order or a Nontrivial Factor of NN
@@
9:   Return $a$ as a nontrivial factor of NN or, if $a=N$, “NN is prime”.
@@
16:     Return $\gcd(N,a^{m_{i}/q}-1)$ as a nontrivial factor of NN
@@
20:     Return $\gcd(N,M_{i+1})$ as a nontrivial factor of NN
@@
By applying 2.6, the first while loop in step 6 takes $O(\sqrt{M_{i}})$ iterations (rather than the $O(M_{i})$ given by [Hit18]). This allows us to modify the condition in step 18 to $M_{i}\geq D$ instead of $M_{i}\geq\sqrt{D}$ without impacting the overall running time.
@@
By applying 2.6 with $\ell\coloneqq D$, we observe that the while loop in step 6 executes at most $O(\sqrt{D})$ iterations. Note that the condition that $N$ has a prime factor $p$ such that $p>2D$ is satisfied, otherwise a factor of $N$ would be found in step 1.
@@
Finally, applying Theorem 3.2 in step 21 with $r\coloneq 1,s\coloneq M_{i+1},\alpha\coloneq 1/6$ is done in time $N^{1/4-1/6+o(1)}$ time, which is $D^{1/2+o(1)}$ since $D\geq N^{1/6}$.